1000 people entered a math contest to see how many questions out of 20 they could answer correctly in 5 minutes. 15% of them answered 12 questions correctly, three-fifths of them answered 15 questions correctly, and the rest answered 18 questions correctly. What was the percent correct that everyone earned in this contest
1
Expert's answer
2019-07-31T09:25:39-0400

1) Number of people:

"1000(0.15)=150"

Proportion of correct questions:

"\frac{12}{20}=0.6"

2) Number of people:


"1000 \cdot \frac{3}{5}=600"


Proportion of correct questions:


"\frac{15}{20}=0.75"

3)Number of people:


"1000-150-600=250"

Proportion of correct questions:


"\frac{18}{20}=0.9"

The proportion correct that everyone earned in this contest:


"\frac{(150)(0.6)+(600)(0.75)+(250)(0.9)}{1000}=0.765"

The percent correct that everyone earned in this contest is 76.5%.